45MM=? CM

2. WHATS LONGER 0.09M 91CM OR 9MM
3.WHAT IS LONGE RFOUR 1 METER PIECES OF WOOD OR SEVEN 60 CM PIECES OF WOOD?
4. A TANK HOLDS 1,000 L OF WATER HAS HOW MANY CUBIC DECIMETERS?
5. 1.68M = ? CM
THANK YOU FOR YOUR HELP
6. IF A SPACE OF 50CM IS NEEDED FOR 1 PERSON TO SIT HOW MANY PEOPLE CAN FIT ON A BENCH 1 DEKAMETER LONG?

Answer by sarah_adam(201)

a) 10 millimeters = 1 centimeter
45 mm = 45/10 = 4.5 cm
b)lets convert everything to centimeters
1 m = 100 cm
0.09 m = 0.09*100 = 9 cm
91 cm
9 mm = 9/10 = 0.9cm
therfore 91 cm is the greatest among three
c)four one meter pieces of wood 4 * 1 = 4m = 400 cm
seven 60 cm pieces of wood = 7*60 = 420 cm
seven 60 cm pieces are larger than 4 one meter pieces of wood
d)1000 cu. centimeters = 1 cu. decimeter
1000 litres = 1000 cubic decimeter
e)10 meters = 1 dekameter
50 cm is required for one person to sit
that is 50/100 = 0.5 m is required for one person
therfore number of persons that can sit on 10 meters long bench are
10/0.5 = 20
20 persons can sit on bench of 1 dekameter long
f)1m = 100 cm
1.68 m = 1.68*100 = 168cm